OpenStudy (anonymous):

what is 16 square root ?

OpenStudy (anonymous):

4

OpenStudy (mattyboyy):

4

OpenStudy (anonymous):

4

sammixboo (sammixboo):

Please do not give direct answers =( They are against the CoC, but you can find the square root of somrthing by finding the number it multiplies itself to get that number

OpenStudy (anonymous):

dont just give me ans. tell me what is square root

sammixboo (sammixboo):

Say you are trying to find the square root of 25. What number do you multiply itself to get 25? 5. Because 5 x 5 = 25

OpenStudy (anonymous):

i dont inderstand

OpenStudy (anonymous):

\[\sqrt{16}=\pm ?\]

OpenStudy (anonymous):

means 16 square root = x*x

sammixboo (sammixboo):

Well, here is a better description. Weird to describe for me http://www.mathsisfun.com/square-root.html

sammixboo (sammixboo):

Correct

OpenStudy (anonymous):

the answer is 4

sammixboo (sammixboo):

You multiply 4 by itself to get 16, right?

OpenStudy (anonymous):

so if x is 4 then square root of 16 is = 4 x 4

OpenStudy (anonymous):

oh thanks a lot

OpenStudy (anonymous):

MEOWWWWWWWWWWW

sammixboo (sammixboo):

It's just 4 =P

sammixboo (sammixboo):

If you have any more problems, tag users by typing @ and their username, like @sammixboo And if you have any concerns about this site Private Message me anytime I am online =)

OpenStudy (anonymous):

ISNT IT 21?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

square root NOUN a number that produces a specified quantity when multiplied by itself: "7 is a square root of 49"

OpenStudy (anonymous):

what about -4*-4=? and 4*4=?

OpenStudy (anonymous):

-4*-4=16 4*4=16

OpenStudy (anonymous):

what do u mena